# Active Preference Learning for Ordering Items In- and Out-of-sample

### v1

@@ C.4 BALD
Where the posterior is calculated as in Appendix C.2 and $H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ is approximated as in Appendix C.4.1.
C.4.1 Deriving the BALD sampling criterion
@@
Here $\Pr(y\mid z_{i,j},D_{t})$ is the predictive distribution for our Bayesian logistic regression model. As covered in Bishop and Nasrabadi (2006), this expectation cannot be evaluated analytically but can be approximated using the probit function $\Phi$;
@@
where $C=\sqrt{4\ln{2}}$. Finally, we arrive at an estimation of the objective function which we wish to maximize:
$H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]\approx h\left(\sigma\left(\frac{\theta_{t}^{\top}z_{i,j}}{\sqrt{1+\frac{\pi}{8}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\right)-\frac{C}{\sqrt{||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2}}}\exp{\left(-\frac{(\theta_{t}^{\top}z_{i,j})^{2}}{(||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2})}\right)}$
Appendix D Proofs of Lemma 1 and Theorem 1
D.1 Proof of Lemma 1
@@ D.2 Proof of Theorem 1
We let $i\succ j$ denote that $i$ is preferred to $j$. W.l.o.g assume $1\succ 2\succ...\succ n$. The key observation is that the assumption $\sigma(z_{ij}^{\top}\theta_{*})-0.5>\Delta_{ij}>\Delta_{*},\forall i,j$ implies that for any $i$ and $j$ such that $i<j$ it holds that
@@
Appendix E Comparison with regret minimization
@@
Appendix F Experiment details
F.1 Ground truth on PhotoAge
While the IMDB dataset used in the PhotoAge task contains an associated age with every image, these were not used. There were two reasons for this. First, some of these ages were incorrectly assigned (often in photos containing multiple people that had been cropped to one individual), causing some noise. Second, and most importantly, we are interested in the subjective assessment of how old someone appears to be, which does not have to be consistent with the true age. The consequence of this is that if you keep collecting comparisons, there is no guarantee that the ordering will ever converge to reflect the true age. The rationale behind using the Borda Score instead is that the simple algorithm does not make any assumptions regarding transitivity Nihar B. Shah (2017), thus not favoring any of the sampling strategies.
F.2 Implementation Details PhotoAge
@@
To increase computational efficiency, the hybrid algorithms did not evaluate all 100 000similar-toabsent100000$\sim$ 100\ 000∼ 100 000 comparisons at every time step. Instead, a subset of 5 00050005\ 0005 000 comparisons was first sampled, and the highest scoring pair in this set was chosen. This resulted in a large speed up, and no noticeable change in performance during evaluation.
F.3 Additional figures
NormMin Comparison
@@ F.3 Additional figures
To highlight the importance of the first-order term in Lemma 1, we evaluated NormMin on the same X-ray ordering task as in Figure 1(a). The results, shown in Figure 5, indicate that not only does the algorithm perform worse than GURO, but is seemingly also outperformed by a uniform sampling strategy.
Figure 5: X-RayAge. NormMin included in the experiment shown in Figure. 1(a)
@@
As discussed in Section 6.2, the performance of our contextual approach will depend on the quality of the representations. To underscore the practical usefulness of our algorithms, we have performed the same experiment as in Figure 4, but this time the model used to extract image features was untrained (i.e. the weights were random). As to be expected, the results, shown in Figure 7, demonstrate that the fully contextual algorithms have no real way of ordering the items according to these uninformative features. However, the hybrid BayesGURO algorithm performs similarly to TrueSkill, despite model misspecification. This is promising, since you may not know in advance how informative the extracted features will be for the target ordering task.
Figure 7: PhotoAge. The same experiment as presented in Figure. 4, but the model used for feature extraction is untrained.

### v2

@@ B.5 BALD
Where the posterior is calculated as in Appendix B.2 and $H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ is approximated as in Appendix B.5.1.
B.5.1 Deriving the BALD sampling criterion
@@
Here $\Pr(y\mid z_{i,j},D_{t})$ is the predictive distribution for our Bayesian logistic regression model. As covered in Bishop and Nasrabadi (2006, Chapter 4), this expectation cannot be evaluated analytically but can be approximated using the probit function $\Phi$;
@@
where $C=\sqrt{4\ln{2}}$. Finally, we arrive at an estimation of the objective function we wish to maximize:
$H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]\approx$ $\ h\left(\sigma\left(\frac{\theta_{t}^{\top}z_{i,j}}{\sqrt{1+\frac{\pi}{8}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\right)$
$-\frac{C}{\sqrt{||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2}}}\exp{\left(-\frac{(\theta_{t}^{\top}z_{i,j})^{2}}{(||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2})}\right)}$
Appendix C Proofs of Lemma 1 and Theorem 1
C.1 Proof of Lemma 1
@@ C.2 Proof of Theorem 1
We let $i\succ j$ denote that $i$ is preferred to $j$. W.l.o.g assume $1\succ 2\succ...\succ n$. The key observation is that for any $i$ and $j$ such that $i<j$ it holds that
@@
C.3 Extensions of current theory
Regularized estimators.
In our analysis in Section 4, we have assumed that $\theta_{T}$ is the maximum likelihood estimate and that $\mathbf{H}(\theta_{T})$ has full rank. This can be relaxed by considering $\ell_{2}$ (Ridge) regularization where $\theta_{\lambda_{0},T}$ is the optimum of the regularized log-likelihood with regularization $\lambda_{0}\mathbf{I}$ and $\mathbf{H}_{\lambda_{0}}(\theta_{\lambda_{0},T})=\sum_{s=1}^{T}\dot{\sigma}(z_{s}^{\top}\theta_{\lambda_{0},T})z_{s}z_{s}^{\top}+\lambda_{0}\mathbf{I}$. The same machinery used to prove Lemma 1 (Filippi et al., 2010; Faury et al., 2020) can be applied to this regularized version with small changes to the final bound.
Generalized linear models.
It is also possible to derive similar results for generalized linear models with other link functions, $\mu(z_{ij}^{\top}\theta_{*})$, by using the general inequality $\mathbf{H}(\theta)\geq\kappa^{-1}\mathbf{V}$ with $\mathbf{V}=\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ and $\kappa\geq 1/\min_{z_{ij}}\dot{\mu}(z_{ij}^{\top}\theta_{*})$. We conjecture that this will yield a scaling of $\sim\exp(-\Delta^{2}T/\kappa)$ where, unfortunately, $\kappa$ might be very large. For a more thorough discussion on the dependence on $\kappa$ in generalized linear bandits, see Lattimore and Szepesvári (2020, Chapter 19).
Lower and algorithm-specific upper bounds on the ordering error.
A worst-case lower bound on the ordering error can be constructed in the fixed-confidence setting, where the goal is to minimize the number of comparisons until a correct ordering is found with a given confidence, by following Garivier and Kaufmann (2016). This involves defining the set of alternative models $\mathrm{Alt}(\theta_{*})$ which differs from $\theta_{*}$ in their induced ordering of $\mathcal{I}$. The bound is then constructed by optimizing the frequency of comparisons of each pair of items so that such alternative models are distinguished as much as possible from the true parameter. We have left this result out of the paper as we find it uninformative in the regime when the number of comparisons is small, (see Simchowitz et al. (2017) for a discussion on the limitations of these asymptotic results in the standard bandit setting). Constructing a lower bound for our fixed-budget setting, of learning as good an ordering as possible with a fixed number of comparisons, is much more challenging. The fixed-confidence result yields a bound for the fixed-budget case (Garivier and Kaufmann, 2016), but constructing either a tight lower bound or a tight algorithm-specific upper bound is an open problem (Fang, 2022).
Appendix D Comparison with regret minimization
@@
Appendix E Experiment details
For BayesGURO and BALD, the posterior $p(\theta\mid D_{t})$ is estimated using the Laplace approximation as described in Bishop and Nasrabadi (2006, Chapter 4). With this approximation, the covariance matrix is the same as the inverse of the Hessian of the log-likelihood. For both methods, the priors $\theta_{B,0}=\mathbf{0}^{d}$ and $\mathbf{H}_{B,0}^{-1}=I_{d}$ were used, and sequential updates were performed every iteration. The sample criterion for BALD under a logistic model is given in Appendix B.5.1. For BayesGURO, 50 posterior samples were used to estimate $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ for every $z_{ij}$. The hybrid algorithms follow the same structure with the added constraint that each per-item parameter $\zeta_{i}$ is independent of other parameters. This allows for efficient updates of $\mathbf{H}_{B,t}^{-1}$ by using sparsity in the covariance.
GURO, CoLSTIM, and Uniform use LogisticRegression from Scikit-learn (Pedregosa et al., 2011) with default Ridge regularization ($C=1$) and the lbfgs optimizer. The former two updates $\theta_{t}$ every iteration using the full history, $D_{t}$ in all experiments except for IMDB-WIKI-SbS, where GURO updates $\theta_{t}$ every 25th iteration. This caused no noticeable change in performance as GURO still updates $\mathbf{H}^{-1}_{t}$ every iteration using the Sherman-Morrison formula. Note that when using the Sherman-Morrison formula in practice, you only get an estimate of $\mathbf{H}^{-1}_{t}(\theta_{t})$ since previous versions have been calculated using older estimates of $\theta$. This method for approximating the inverse hessian is covered in Bishop and Nasrabadi (2006, Chapter 5) and when we compared it to calculating $\mathbf{H}^{-1}_{t}(\theta_{t})$ from scratch every iteration we observed that the methods performed equally. The design matrix for CoLSTIM is updated as in Bengs et al. (2022): the confidence width $c_{1}$ was chosen to be $\sqrt{d\log(T)}$, and the perturbed values were generated using the standard Gumbel distribution.
@@ Appendix E Experiment details
To increase computational efficiency for the large IMDB-WIKI-SbS dataset, the hybrid algorithms did not evaluate all 100 000similar-toabsent100000$\sim$ 100\ 000∼ 100 000 comparisons at every time step. Instead, a subset of 5 00050005\ 0005 000 comparisons was first sampled, and the highest-scoring pair in this set was chosen. This resulted in a large speed-up and no noticeable change in performance during evaluation.
E.1 Datasets
ImageClarity
Data available at https://dbgroup.cs.tsinghua.edu.cn/ligl/crowdtopk. This dataset contained differently distorted versions of the same image. To extract relevant features, we used a ResNet34 model (He et al., 2016) that had been pre-trained on Imagenet (Deng et al., 2009). After PCA projection feature dimensionality was reduced to $d=63$. The dataset consisted of $100$ images and 27 7302773027\ 73027 730 comparisons. Since the type of distortion is the same for all images, the dataset has a true ordering with regards to the strength of the distortion applied.
WiscAdds
Data available at https://dataverse.harvard.edu/dataset.xhtml?persistentId= doi:10.7910/DVN/0ZRGEE (license: CC0 1.0). The WiscAdds dataset, containing $935$ political texts, has been extended with 9 52895289\ 5289 528 pairwise comparisons by Carlson and Montgomery (2017). In comparisons, annotators have stated which of two texts has a more negative tone toward a political opponent. To extract general features from the text, sentences were embedded using the pre-trained all-mpnet-base-v2 model from the Sentence-Transformers library (Reimers and Gurevych, 2019). After applying PCA to the sentence embeddings, each embedding had a dimensionality of $d=162$.
IMDB-WIKI-SbS
Data available at https://github.com/Toloka/IMDB-WIKI-SbS (license: CC BY). IMDB-WIKI-SbS consists of close-up images of actors of different ages. For each comparison, the label corresponds to which of two people appears older. The complete dataset consists of 9 15091509\ 1509 150 images and 250 249250249250\ 249250 249 comparisons, but images that were grayscale or had a resolution lower than $160\times 160$ were removed, resulting in 6 072 images and 110 349 comparisons. We extract features from each image using the Inception-ResNet implemented in FaceNet (Schroff et al., 2015) followed by PCA, resulting in $d=75$ features per image.
E.2 Additional figures
X-RayAge
@@ E.2 Additional figures
To highlight the importance of the first-order term in Lemma 1, we evaluated NormMin on the same X-ray ordering task as in Figure 1(a). The results, shown in Figure 7(a), indicate that not only does the algorithm perform worse than GURO, but is seemingly also outperformed by a uniform sampling strategy. Furthermore, for completeness, we include Figure 7(b) which shows the in-sample error, $R_{I_{D}}$, during the generalization experiment.
(a) X-RayAge. NormMin included in the experiment shown in Figure 1(a).
(b) X-RayAge. The in-sample error $R_{I_{D}}$ for the generalization experiment performed in Figure 1(b).
Figure 7: Additional figures from the X-RayAge experiment.
@@
As discussed in Section 1, the performance of our contextual approach will depend on the quality of the representations. To underscore the practical usefulness of our algorithms, we have performed the same experiment as in Figure 6, but this time the model used to extract image features was untrained (i.e., the weights were random). As to be expected, the results, shown in Figure 9, demonstrate that the fully contextual algorithms have no real way of ordering the items according to these uninformative features. However, GURO Hybrid performs similarly to TrueSkill, despite model misspecification. This is promising, since you may not know in advance how informative the extracted features will be for the target ordering task.
Figure 9: IMDB-WIKI-SbS. The same experiment as presented in Figure. 6, but the model used for feature extraction is untrained.
ImageClarity ground truth
The ImageClarity dataset consists of multiple versions of the same image, with the same distortion applied to it to varying degrees. Due to this artificial construction, the pairwise comparisons should, given enough samples, reflect the magnitudes of the applied distortions. In Figure 10 we perform the same experiment as in Figure 6, but instead of evaluating on a holdout comparison set, we measure the distance to the ground-truth ordering. The overall results are very similar, although we do see a slight increase in the performance of contextual algorithms compared to the non-contextual TrueSkill.
Figure 10: ImageClarity. Same experiment as in Figure 6, but now measuring the distance to the ground-truth ordering. Averaged over 25 seeds along with the 1-sigma error region.
Ground truth ordering using the Bradley-Terry model
An alternate approach to evaluate ordering quality is to estimate a ”ground-truth” ordering by applying the popular Bradley-Terry (BT) model (Bradley and Terry, 1952) to all available comparisons. We used the CrowdKit library (Ustalov et al., 2024) to find the MLE scores for each item and ordered the elements accordingly. In Figure 14 we run the same experiments as in Figure 6, but instead measure the distance to the constructed BT ordering. The overall trends remain, but for (14) and (14) there is a slight shift for the later iterations. More specifically we see non-contextual TrueSkill eventually overtaking the contextual algorithms.
{subcaptionblock}
[t]0.49 Refer to caption {subcaptionblock}[t]0.49 Refer to caption
Figure 11: ImageClarity. $n=100$, $d=63$.
Figure 12: WiscAds. $n=935$, $d=162$.
{subcaptionblock}
[t]0.49 Refer to caption
[…]
Figure 14: The same experiment as presented in Figure. 6, but we instead measure the distance to a ground-truth estimated using all available comparisons.
The issue is that algorithms with orderings closer to the maximum likelihood estimate of the BT model will be favored. To exemplify this we use the ImageClarity dataset since it contains the largest number of comparisons relative to the number of items. We sample 1 00010001\ 0001 000 comparisons and let this be the collection that is available to the algorithms. We further construct two target orderings, one from the BT estimate using the sampled subset of comparisons, and a second, more probable ordering, from the BT estimate using all 27 7302773027\ 73027 730 available comparisons. Figure 15 shows the distance between the GURO and TS algorithms and the different target orderings, where dashed lines indicate the distance to the ordering generated using the full list of comparisons. If we only look at the distance to the ordering produced using our subset of comparisons, TrueSkill seemingly outperforms GURO after about $350$ comparisons. However, if we instead measure the distance to the more probable ordering, we see that GURO converges toward a lower distance. Note that these are the same orderings, evaluated against different targets. This is likely the effect we observe in Figure 14 and 14, but not in Figure 14 as a result of the high amount of comparisons available to us.
Figure 15: ImageClarity. The same experiment as presented in Figure 6, but we instead measure the distance to target orderings that correspond to the maximum likelihood estimate of the BT model using different numbers of comparisons. The dashed lines show the distance to the BT estimate using all 27 7302773027\ 73027 730 comparisons.